OpenStudy (anonymous):

heelp i will medal and faan :) Solve for x: one over eight (8x + 15) = 24

OpenStudy (anonymous):

x= 22.125 i checked this time

TheSmartOne (thesmartone):

\[\frac{ 1 }{ 8 }(8x+15)=24\] \[(8x+15)=24\times8\] \[8x+15=192\] \[8x=192-15\] \[8x=177\] \[x=\frac{ 177 }{ 8 }\] \[x=22.125\]